1. The heat capacity of a bottle of water is 2100 J°C-1. What is the amount of heat required to heat the water from 30°C to 50°C?
2.1 kg of substance X of specific heat capacity 2 kJkg-1°C-1 is heated from 30°C to 90°C. Assuming no heat loss, the heat required is?
1
Expert's answer
2020-05-07T10:00:05-0400

1)


"Q=C\Delta T=(2100)(50-30)=42000\ J"

2)


"Q=mc\Delta T=(2.1)(2000)(90-30)=250000\ J"